OpenStudy (morganchavez):

ALGEBRA 2, UNIT 3, LESSON 5: Graphing Quadratics ALTERNATE ASSIGNMENT Did you know parabolas are all around you in the real world!? If you’ve traveled to the Midwest, you may have seen this one before: The Gateway Arch in St. Louis, Missouri is a 630-foot high monument that was first designed in 1947. Construction of the stainless steel arch lasted many years, and the arch was not completed until October 1965. It was created as a monument to the westward expansion of the United States, and to this day it is the tallest man-made monument in America, as well as the world’s tallest arch

OpenStudy (morganchavez):

Use the following function to complete #1-3 below f(x)=2(x-6)^2+4 1) Using complete sentences, explain how to convert f(x) into standard form. Give the final standard form equation in your answer. 2) Graph f(x) by hand or with technology. Include a photo or screenshot of the parabola. 3) Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

hey XD

OpenStudy (morganchavez):

hi lol im still stuff on this problem for a week :(

OpenStudy (morganchavez):

@RANE @mathgeek27 @mathmath333 @rwestfall_ca @MakaylaTracy @walters

OpenStudy (anonymous):

Its ok were here to help YOU

OpenStudy (anonymous):

i might be able to help but then again im only a freshman :(

OpenStudy (morganchavez):

thanks i really need it pple gave me the answers but never told me how they got it n what i need is lik to say how i got the answer in complete sentences and so im so lost

OpenStudy (morganchavez):

i understand @J-bird

OpenStudy (anonymous):

\[ f(x)=2(x-6)^2+4 \] 1) Using complete sentences, explain how to convert f(x) into standard form. Multiply out, combine like terms

OpenStudy (anonymous):

i could do the problem and graph it and all but i dont think i could explain :P

OpenStudy (anonymous):

kind of a short sentence, but that is all there is to it \[f(x)=2(x-6)^2+4=2(x-6)(x-6)+4=...\]

OpenStudy (morganchavez):

the standard form of the porabola is y=2x^2-27x+76 which is #1

OpenStudy (morganchavez):

right ?

OpenStudy (anonymous):

Good Job @satellite73

OpenStudy (anonymous):

2) Graph f(x) by hand or with technology. Include a photo or screenshot of the parabola. well then lets use technology ! http://www.wolframalpha.com/input/?i=2%28x-6%29^2%2B4

OpenStudy (morganchavez):

ok :)

OpenStudy (anonymous):

btw technology also multiplied out for you, you see it is \[f(x)=2 x^2-24 x+76\]

OpenStudy (morganchavez):

wait but which one would it be ?

OpenStudy (anonymous):

which is not what you had

OpenStudy (morganchavez):

ooo :( wow thank you if u didnt catch thst i would hsve lost major points and i cant aaford that

OpenStudy (anonymous):

the pictures are both the same, only the scale is different i would go with the top one though

OpenStudy (morganchavez):

ok :)

OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. average rate of change is f of seven minus f of four divided by seven minus four jeez the sentence is kind of silly, it is \[\frac{f(7)-f(4)}{7-4}\]

OpenStudy (anonymous):

tell FLVS this "complete sentence" nonsense is written by morons you can compute those numbers i am sure good luck

OpenStudy (morganchavez):

? so 3 divided by 3 lol

OpenStudy (morganchavez):

i will lol half the work they give me make no sense or have a million typos

OpenStudy (anonymous):

i didn't check it \(7-4=3\) for sure you want me to check the numerator also?

OpenStudy (anonymous):

hey i'm FLVS :( ;*(

OpenStudy (morganchavez):

yes please

OpenStudy (anonymous):

ok it is wrong

OpenStudy (morganchavez):

? what do u mean

OpenStudy (just_one_last_goodbye):

@cuteandsexy thats great please post your question in your post and no here thats spam ^_^ thank you for complying

OpenStudy (anonymous):

\[f(7)=2(7-6)^2+4=2\times 1+4=6\]

OpenStudy (anonymous):

\[f(4)=2(4-6)^2+4=2\times 4+4=12\]

OpenStudy (anonymous):

\[f(6)-f(4)=6-12=-6\]

OpenStudy (anonymous):

and \[\frac{f(7)-f(4)}{7-4}=\frac{-6}{3}=-2\]

OpenStudy (morganchavez):

0_0 ? wait how would it equal -6 it makes no sense i thought it was 3/3?

OpenStudy (anonymous):

i don't know another way to say it compute \[f(7)\] i get \[f(7)=6\] compute \[f(4)\] i get \[f(4)=12\]

OpenStudy (anonymous):

then compute \[\frac{f(7)-f(4)}{7-4}=\frac{6-12}{3}=\frac{-6}{3}=-2\]

OpenStudy (morganchavez):

are u using wolf of alpha if so it gives wrong answers sometimes try mathway ...i dont show u the work but it does the answer

OpenStudy (anonymous):

i did it by hand

OpenStudy (morganchavez):

nope never mind that is right wow math has many secrets

OpenStudy (anonymous):

do you know how to compute \(f(7)\) ?

OpenStudy (morganchavez):

no i put it in the math site and yup u were right im so sorry i wasnt telling u u were wrong i jus thought it was odd bc it didnt make sense at first lol

OpenStudy (morganchavez):

@phi is all this right ?

OpenStudy (phi):

You could make it your goal to learn how to figure out f(7) \[ f(x)=2(x-6)^2+4 \] f(x) is a "short name" for the formula the (x) part is to remind us the formula uses "x" f(7) is short for: find the "x" in the formula. Erase it, and put in its place the number 7 can you do that ?

OpenStudy (morganchavez):

I JUST WANNA MAKE DOULE SURE BEFORE I WRITE IT ALL OUT AND TRUN IT IN im lik almost a whole week overdue

OpenStudy (morganchavez):

yes i think yes

OpenStudy (morganchavez):

yes i can sorry yes i can replace all x's with the number 7 :)

OpenStudy (phi):

write down the formula with x replaced with 7

OpenStudy (phi):

you should get f(x)=2(x-6)^2+4 f(7) = 2(7-6)^2 + 4

OpenStudy (phi):

Do you remember your "order of operations" ? to simplify 2(7-6)^2 + 4 what do we do first ?

OpenStudy (morganchavez):

pemdas?

OpenStudy (morganchavez):

paranthesis

OpenStudy (phi):

yes, and what do you get ? i.e. 7-6

OpenStudy (morganchavez):

so 2(1)^2+4 then we do exponents

OpenStudy (phi):

yes

OpenStudy (morganchavez):

so then we get 2(1)+4 then we multiply

OpenStudy (morganchavez):

so we get 2+4 and then we dont have any division to do so we go to addition and get 6 as our answer right ?

OpenStudy (morganchavez):

what do we do next ?

OpenStudy (phi):

exactly f(7) = 6 that means when x is 7, f(7) is 6 or (7,6) is a point on the curve.

OpenStudy (phi):

average rate of change for f(x) from x = 4 to x = 7. the "rate of change" is change in f(x) divided by change in x in other words we have to figure out f(7) - f(4) (that is how much f(x) changes between x=4 and x=7) so we need to find f(4) any idea how ?

OpenStudy (morganchavez):

O_O wow i didnt know that omg thank you can u help me with more like is my problem rightt ? the one that i did with sattelitte ? and now i know omg thanks my teacher in flvs doesnt teach sqwat

OpenStudy (morganchavez):

yes hang on n ill right it out

OpenStudy (morganchavez):

2(4-6)^2+4 right ?

OpenStudy (phi):

yes

OpenStudy (morganchavez):

so then i do paranthisis -2 so i get 2(-2)^2+4

OpenStudy (phi):

yes

OpenStudy (morganchavez):

then exponents 4 so 2(4)+4 then i multiply and i get 8+4 and since there is no division i do addition so i get 12 as my answer right ?

OpenStudy (phi):

yes, so when x is 4, f(x) is 12 (4,12) is also a point on the curve. the vertex (which we can get from the equation) is (6,4) if we plot those 3 points, it looks like this

OpenStudy (phi):

now we can find the difference between f(7) and f(4) f(7) - f(4) is ?

OpenStudy (morganchavez):

wait im sorry so interupt but why is there 3 points?

OpenStudy (phi):

Oh, I know (and probably they want you to know) that when you have an equation that looks like this: a (x - h)^2 + k that the vertex (the top or bottom of the parabola) is at (h,k) if we match that formula with 2(x-6)^2+4 we see the h matches with 6 and k with 4 the vertex is at (h,k) , which is (6,4) I did that to help you plot this curve.

OpenStudy (phi):

meanwhile, we now want to find f(7) - f(4) find the numbers for f(7) and f(4) which you figured out. and do the subtraction can you do that ?

OpenStudy (morganchavez):

yes well i know f(7) is 6

OpenStudy (phi):

yes

OpenStudy (morganchavez):

f(4) is 12

OpenStudy (phi):

yes

OpenStudy (morganchavez):

so then i do 6-12

OpenStudy (phi):

yes

OpenStudy (morganchavez):

and i get -6

OpenStudy (phi):

that says from point A to point C (in the graph I posted), we moved down 6 steps (from 12 down 6 steps to get to 6) the average rate of change for f(x) from x = 4 to x = 7. is the change in f(x) divided by the change in x the change in x is 7-4 = 3 now we do change in f(x) divided by change in x i.e. (f(7)-f(4) ) / (7-4)

OpenStudy (phi):

btw, if we drew a line from point A to point C, we are finding the slope of that line slope is (change in y)/(change in x) and we usually match f(x) with y when we plot

OpenStudy (morganchavez):

6-12/6-12 right ?

OpenStudy (morganchavez):

or is it 6-12/ (7-4)?

OpenStudy (phi):

the second. change in f(x) divided by change in x or, if you know how to find the slope between 2 points, it is the slope between (4,12) and (7,6)

OpenStudy (morganchavez):

wait so 4-12 divided by 7-6 so 4-12/7-6

OpenStudy (phi):

and it would be good to write it as (6-12)/ (7-4) because if we followed order of operations 6 - 12/(7-4) would do the wrong thing.

OpenStudy (morganchavez):

so -8/1

OpenStudy (morganchavez):

ooooooo bc its y1-y2/x1-x2 right ?

OpenStudy (phi):

I think I confused you with the slope idea. it is the difference in the y values (the 2nd number) divided by the difference in the x values.

OpenStudy (phi):

*** 4-12 divided by 7-6 that is x-y divided by x-y which is not what we want

OpenStudy (morganchavez):

oooo ok sorry my bad lol i had my formulas for slope wrong my bad

OpenStudy (morganchavez):

si its x1-y1/ y2-x2 right ?

OpenStudy (phi):

*** y1-y2/x1-x2 right ? this is correct. But I would use parens (y1-y2)/(x1-x2)

OpenStudy (phi):

just be careful and think out each step. pick out y1 pick out y2 write it down then do the same for the x's